OpenStudy (anonymous):

Find the missing value to the nearest hundredth.x = missing value sin30 = (5)/(x) tan ____ = 23 cos 40 = (x)/(8)

OpenStudy (mertsj):

\[\sin 30=\frac{5}{10}\]

OpenStudy (mertsj):

\[\tan 87.5=23\]

OpenStudy (mertsj):

\[\cos 40=\frac{6.128}{8}\]

OpenStudy (anonymous):

thank you very much!

OpenStudy (mertsj):

yw. But I did not do them to the nearest hundredth. first one would be .50 second one: 87.51 last one: 6.13